OpenStudy (anonymous):

Which choice describes the translation represented by the translation rule (x,y) -->(x+4, y-1)? (1 point) 4 units to the right and 1 unit down 1 unit to the right and 4 units down 4 units to the left and 1 unit up 4 units to the left and 1 unit down

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x to x+4 means you're taking some x coordinate and turning it into x+4

jimthompson5910 (jim_thompson5910):

visually, you're translating the point 4 units to the right

jimthompson5910 (jim_thompson5910):

y to y-1 means what?

OpenStudy (anonymous):

thanks. I already figured it out though. xP

jimthompson5910 (jim_thompson5910):

ok great, glad you were able to figure it out

OpenStudy (anonymous):

Can you help with this? The hexagon GIKMPR and FJN are regular. The dashed line segments form 30° angles. Find the angle of rotation about O that maps JK to FG. 90° 240° 150° 120°

jimthompson5910 (jim_thompson5910):

one sec while I draw

OpenStudy (anonymous):

K

jimthompson5910 (jim_thompson5910):

how many angles have I highlighted?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

each angle is 30 degrees, so the answer is 5*30 = 150

OpenStudy (anonymous):

What about this one? The hexagon GIKMPR and FJN are regular. The dashed line segments form 30° angles. Find the angle of rotation about O that maps P to G. (1 point) 240° 120° 210° 270°

OpenStudy (anonymous):

the picture is the same

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok drawing it out

jimthompson5910 (jim_thompson5910):

how many angles did I mark below?

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

so 8*30 = 240 is your answer

OpenStudy (anonymous):

why times 30?

OpenStudy (anonymous):

and I really don't get this. Find the degree of rotation about the spinner center that maps label i to label g. (1 point) 72° 108° 36° 288°

jimthompson5910 (jim_thompson5910):

because each angle I highlighted is 30 degrees

jimthompson5910 (jim_thompson5910):

the other one, you're going 2 spots over each angle is 36 degrees, so you're going a total of 2*36 = ??? degrees

OpenStudy (anonymous):

72

jimthompson5910 (jim_thompson5910):

yep